OpenStudy (anonymous):

The levels of mercury in two different bodies of water are rising. In one body of water the initial measure of mercury is 0.05 parts per billion (ppb) and is rising at a rate of 0.1 ppb each year. In the second body of water the initial measure is 0.12 ppb and the rate of increase is 0.06 ppb each year. After how many years will the amount of mercury be the same? a. 0.5 years b. 1.25 years c. 1.75 years d. 2.5 years

OpenStudy (anonymous):

@dumbo

OpenStudy (anonymous):

c is my best bet

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

this is really confusing im sorry

OpenStudy (anonymous):

Your fine.

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

let after x year it will be same

OpenStudy (gorv):

0.05+0.1*x=0.12+0.06*x subtract 0.05 from both side

OpenStudy (gorv):

0.1x=0.12-0.05+0.06x 0.1x=0.07+0.06x subtract 0.06x from both side 0.1x-0.06x=0.07 0.04x=0.07 divide both side by o.o4 x=0.07/0.04 x=7/4 year

OpenStudy (gorv):

x= 1 yeah 9 month

OpenStudy (anonymous):

9 months isn't an option..

OpenStudy (gorv):

loll 7/4 year

OpenStudy (gorv):

what are the option

OpenStudy (gorv):

1.75 year

OpenStudy (gorv):

1 3/4

OpenStudy (gorv):

\[\frac{ 7 }{ 4 }=1.75=1 year 9 month ,,,,,, all are same\]